OpenStudy (anonymous):

Please help me... Graphs anyone? I have an equation, thats it... :) Given f(x) and g(x) = f(x) + k, look at the graph below and determine the value of k. k = ______

OpenStudy (anonymous):

so 1/3x+3=1/3x−2+k

OpenStudy (anonymous):

OR g(x)-f(x)=k

OpenStudy (loser66):



OpenStudy (loser66):